OpenStudy (amistre64):

Prove: [m^2 = n^2] iff [(m = n) or (m = -n)]. ugh ....

OpenStudy (amistre64):

to prov a biconditional you split it into its conditional parts and prove

OpenStudy (amistre64):

p <-> q p -> q AND q -> p

OpenStudy (anonymous):

?

OpenStudy (anonymous):

isn't it obvious.

myininaya (myininaya):

m^2-n^2=(m-n)(m+n)=0 => m=n or m=-n

myininaya (myininaya):

that is p->q

OpenStudy (amistre64):

lol ... thats what I thought; why prove the obvious :)

OpenStudy (amistre64):

ahh, diff of squares ..... i keep trying tog o to basic with this stuff and prove that counting exists .....

myininaya (myininaya):

you got q->p?

OpenStudy (amistre64):

lol .... im still working on if (amistre64) then (exist)

OpenStudy (amistre64):

im thinking contradiction or contraposition in that one

OpenStudy (amistre64):

(m=n or m=-n) -> (m^2 = n^2) If (m^2 ≠ n^2), then (m≠n and m≠-n) then follow as before

OpenStudy (amistre64):

unless there is something glaringly obvious that I am missing

OpenStudy (amistre64):

proof by contradiction would be in line with the previous proof tho and might be simpler

OpenStudy (amistre64):

if (m=n or m=-n) then (m^2 \(\ne\) n^2)

OpenStudy (anonymous):

i hate proofs by contradiction

myininaya (myininaya):

why can't you just say if m=n, then m^2=n^2 and why can't you just say if m=-n, then m^2=(-n)^2=n^2 is that just too easy?

OpenStudy (anonymous):

if \[m^2=n^2\iff m^2-n^2=0\iff (m+n)(m-n)=0\] and then it is an openstudy problem

myininaya (myininaya):

satellite cheated off me

OpenStudy (amistre64):

i think we have to do the books method of splitting the hairs

OpenStudy (amistre64):

lets ban him ;)

OpenStudy (anonymous):

this is one direction for sure. i just woke up no banning me! i can't even see straight yet

myininaya (myininaya):

i'm gonna type 525325235 for the amout of hours of suspension

OpenStudy (anonymous):

you would be doing me a favor. i have at least 4 hours of work to do before i leave the house

myininaya (myininaya):

lol

OpenStudy (anonymous):

other direction is trivial. if \[m=n \text { then } m^2=n^2\] and if \[m=-n \text { then } m^2=(-n)^2 = n^2\]

myininaya (myininaya):

omg cheater

OpenStudy (anonymous):

did you write that? ok i cheated, but you have to admit it is the obvious thing to write. what else?

myininaya (myininaya):

i didn't put in pretty latex though so i guess i will give you some cool points

OpenStudy (anonymous):

oops that is what you wrote above isn't it? well lets see... you could always take the log , use a some property, and then exponentiate...

OpenStudy (amistre64):

I proofed by contradiction on the 2nd one :)

OpenStudy (anonymous):

if i don't get to work i am going to hate myself by noon. time to get off the computer. @amistre i hate proofs by contradiction. but why now?

myininaya (myininaya):

don't go to work

myininaya (myininaya):

working for free is funner

myininaya (myininaya):

so you can stay with us

OpenStudy (amistre64):

if (m=n or m=-n) then (m^2 ≠ n^2) (m+n)(m-n) = 0 m^2 +mn -mn -n^2 = 0 m^2 - n^2 = 0 m^2 = n^2

OpenStudy (anonymous):

that should annoy your teacher sufficiently

OpenStudy (amistre64):

yay!! my proofs tend to be more socratic :)

myininaya (myininaya):

lol

OpenStudy (anonymous):

having fun?

OpenStudy (amistre64):

always and for never